OpenStudy (anonymous):

What is the solution to the system of equations? 2x – y = 7 y = 2x + 3

OpenStudy (butterflydreamer):

substitute y=2x + 3 into 2x - y = 7 So in your first equation " 2x - y = 7 " , wherever you see "y", you will REPLACE this "y" with 2x + 3 since " y = 2x + 3" :)

OpenStudy (butterflydreamer):